# Backpropagation-free Spiking Neural Networks with the Forward-Forward Algorithm

### v1

@@ 5 Conclusion
This paper presented a novel approach to training spiking neural networks using the Forward-Forward algorithm. Unlike traditional backpropagation, the FF algorithm offers a more biologically plausible and computationally efficient training method relative to BP by utilizing two forward passes instead of the typical forward and backward pass. This approach addresses key limitations of backpropagation, such as the need for global feedback pathways, explicit storage of all intermediate activities, and weight symmetry, all of which are biologically implausible.
Our experimental results demonstrated the effectiveness of the FF algorithm in training SNNs across both spiking (Neuro-MNIST, SHD) and non-spiking (MNIST, Fashion-MNIST, CIFAR-10) datasets. The proposed FF-based SNN model achieved competitive accuracy, often outperforming traditional backpropagation-based SNNs, particularly on MNIST and Fashion-MNIST. Additionally, on complex datasets such as CIFAR-10, our model showed promising performance despite the inherent challenges posed by high-dimensional data in fully connected architectures.
The FF algorithm’s greater adaptability to neuromorphic hardware architectures further underscores its potential for real-world applications, particularly in scenarios requiring low power consumption and on-the-fly learning. Additionally, the algorithm’s inherent advantages, such as local learning and the ability to handle black-box components, make it a versatile tool for future advancements in neuromorphic computing. Future work could explore further optimization of the FF algorithm, its integration with other neuromorphic hardware, and its application to more complex and diverse tasks, paving the way for more robust and scalable spiking neural network solutions.

### v2

@@ 5 Conclusion
This paper presented a novel approach to training spiking neural networks using the Forward-Forward algorithm. Unlike traditional backpropagation, the FF algorithm offers a more biologically plausible training method relative to BP by utilizing two forward passes instead of the typical forward and backward pass. This approach addresses key limitations of backpropagation, such as the need for global feedback pathways, explicit storage of all intermediate activities, and weight symmetry, all of which are biologically implausible. Our experimental results demonstrated the effectiveness of the FF algorithm in training SNNs across some spiking (Neuro-MNIST, SHD) and non-spiking (MNIST, Fashion-MNIST, K-MNIST) datasets. The proposed FF-based SNN model achieved competitive accuracy, often outperforming traditional backpropagation-based SNNs. The compatibility of the Forward-Forward algorithm with spiking-based architectures enhances its suitability for deployment in practical settings, especially those that impose constraints on power efficiency and require rapid, online learning capabilities. Moreover, the algorithm’s inherent advantages—such as layer-wise local learning and its ability to function effectively with black-box components—position it as a promising framework for future advancements in specialized domains, particularly neuromorphic computing. Future research could focus on optimizing the FF algorithm, integrating it with emerging neuromorphic platforms, and applying it to increasingly complex and diverse tasks, thereby enabling the development of more robust and widely applicable spiking neural network solutions.